OpenStudy (lukecrayonz):

y+(6/y)=5 y+(6/y)=5 @Mathematics

OpenStudy (lukecrayonz):



OpenStudy (lukecrayonz):

My online calculator says 3 or 2, but I dont know how they got that.

OpenStudy (anonymous):

ur online calculator is right buddy i feel i sud really take a snap now

OpenStudy (anonymous):

multiply by y yo cancel the denominator

OpenStudy (anonymous):

y^2 -5y + 6 = 0 y^2 -3y - 2y + 6 = 0 y*(y- 3) -2 * (y-3) = 0 (y-2) * (y-3) = 0 y = 2 or, y = 3

OpenStudy (anonymous):

y^2+6=5y y1=2 y2=3

OpenStudy (lukecrayonz):

none of these helped D: please work it out fully